An ant is at one of the corners of cube with sides 10.m what is the mininum distance in meter that the ant has to crawl to reach the farthest corner on the cube

Grade:12th pass

1 Answers

Khimraj
5 years ago

Since it is a ant,so it can move in only 2 -d.Hence the minimum distance is

= \sqrt{(2a)^2 + a^2} = \sqrt{5}a.

Hope it clears............